# Overdense fireworks in GOODS-N: Unveiling a record number of massive dusty star forming galaxies at z5.2 with the N2CLS

### v1

@@ Acknowledgements.
We acknowledge financial support from the Programme National de Cosmologie and Galaxies (PNCG) funded by CNRS/INSU-IN2P3-INP, CEA and CNES, France, and from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (project CONCERTO, grant agreement No 788212). This work is based on observations carried out under project numbers 192-16 with the IRAM 30-m telescope, and projects W21CV, W23CX, and S24CF with NOEMA. IRAM is supported by INSU/CNRS (France), MPG (Germany) and IGN (Spain). This work is also based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. The observations are associated with programs #1181, 1895 and 3577. Some of the data products used here were retrieved from the Dawn JWST Archive (DJA). DJA is an initiative of the Cosmic Dawn Center (DAWN), which is funded by the Danish National Research Foundation under grant DNRF140. We thank Eric F. Jiménez-Andrade for sharing their data products from the 10 GHz Karl G. Jansky Very Large Array survey of GOODS-N. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. GL thanks Joris Witstok and Thomas Henning for enlightening discussions on dust production, and warmly thanks Fengwu Sun for sharing his Python code used to generate Figure 4.References
- Akins et al. (2023) Akins, H. B., Casey, C. M., Allen, N., et al. 2023, ApJ, 956, 61
- Alcalde Pampliega et al. (2019) Alcalde Pampliega, B., Pérez-González, P. G., Barro, G., et al. 2019, ApJ, 876, 135
- Arrabal Haro et al. (2018) Arrabal Haro, P., Rodríguez Espinosa, J. M., Muñoz-Tuñón, C., et al. 2018, MNRAS, 478, 3740
- Barger et al. (2022) Barger, A. J., Cowie, L. L., Blair, A. H., & Jones, L. H. 2022, ApJ, 934, 56
- Barro et al. (2019) Barro, G., Pérez-González, P. G., Cava, A., et al. 2019, ApJS, 243, 22
- Barrufet et al. (2023) Barrufet, L., Oesch, P. A., Bouwens, R., et al. 2023, MNRAS, 522, 3926
- Battisti et al. (2020) Battisti, A. J., Cunha, E. d., Shivaei, I., & Calzetti, D. 2020, ApJ, 888, 108
- Behroozi et al. (2019) Behroozi, P., Wechsler, R. H., Hearin, A. P., & Conroy, C. 2019, MNRAS, 488, 3143
- Berta et al. (2025) Berta, S., Lagache, G., Beelen, A., et al. 2025, A&A, 696, A193
- Berta et al. (2013) Berta, S., Lutz, D., Santini, P., et al. 2013, A&A, 551, A100
- Berta et al. (2023) Berta, S., Stanley, F., Ismail, D., et al. 2023, A&A, 678, A28
- Béthermin et al. (2017) Béthermin, M., Wu, H.-Y., Lagache, G., et al. 2017, A&A, 607, A89
- Bing (2022) Bing, L. 2022, PhD thesis, thèse de doctorat dirigée par Lagache, Guilaine Physique et sciences de la matière. Astrophysique et cosmologie Aix-Marseille 2022
- Bing et al. (2023) Bing, L., Béthermin, M., Lagache, G., et al. 2023, A&A, 677, A66
- Bolzonella et al. (2000) Bolzonella, M., Miralles, J. M., & Pelló, R. 2000, A&A, 363, 476
- Boquien et al. (2019) Boquien, M., Burgarella, D., Roehlly, Y., et al. 2019, A&A, 622, A103
- Boylan-Kolchin (2023) Boylan-Kolchin, M. 2023, Nature Astronomy, 7, 731
- Brinch et al. (2024) Brinch, M., Greve, T. R., Sanders, D. B., et al. 2024, MNRAS, 527, 6591
- Brinch et al. (2025) Brinch, M., Jin, S., Gobat, R., et al. 2025, A&A, 694, A218
- Bruzual & Charlot (2003) Bruzual, G. & Charlot, S. 2003, MNRAS, 344, 1000
- Burgarella et al. (2005) Burgarella, D., Buat, V., & Iglesias-Páramo, J. 2005, MNRAS, 360, 1413
- Calvi et al. (2023) Calvi, R., Castignani, G., & Dannerbauer, H. 2023, A&A, 678, A15
- Calvi et al. (2021) Calvi, R., Dannerbauer, H., Arrabal Haro, P., et al. 2021, MNRAS, 502, 4558
- Calzetti et al. (2000) Calzetti, D., Armus, L., Bohlin, R. C., et al. 2000, ApJ, 533, 682
- Carilli & Walter (2013) Carilli, C. L. & Walter, F. 2013, ARA&A, 51, 105
- Carnall et al. (2024) Carnall, A. C., Cullen, F., McLure, R. J., et al. 2024, MNRAS, 534, 325
- Carnall et al. (2018) Carnall, A. C., McLure, R. J., Dunlop, J. S., & Davé, R. 2018, MNRAS, 480, 4379
- Ceverino et al. (2024) Ceverino, D., Nakazato, Y., Yoshida, N., Klessen, R. S., & Glover, S. C. O. 2024, A&A, 689, A244
- Chabrier (2003) Chabrier, G. 2003, ApJ, 586, L133
- Charlot & Fall (2000) Charlot, S. & Fall, S. M. 2000, ApJ, 539, 718
- Chiang et al. (2013) Chiang, Y.-K., Overzier, R., & Gebhardt, K. 2013, ApJ, 779, 127
- Chiang et al. (2017) Chiang, Y.-K., Overzier, R. A., Gebhardt, K., & Henriques, B. 2017, ApJ, 844, L23
- Cowie et al. (2017) Cowie, L. L., Barger, A. J., Hsu, L. Y., et al. 2017, ApJ, 837, 139
- da Cunha et al. (2008) da Cunha, E., Charlot, S., & Elbaz, D. 2008, MNRAS, 388, 1595
- da Cunha et al. (2013) da Cunha, E., Walter, F., Decarli, R., et al. 2013, ApJ, 765, 9
- da Cunha et al. (2015) da Cunha, E., Walter, F., Smail, I. R., et al. 2015, ApJ, 806, 110
- Dekel et al. (2023) Dekel, A., Sarkar, K. C., Birnboim, Y., Mandelker, N., & Li, Z. 2023, MNRAS, 523, 3201
- Di Cesare et al. (2023) Di Cesare, C., Graziani, L., Schneider, R., et al. 2023, MNRAS, 519, 4632
- Draine (2003) Draine, B. T. 2003, ARA&A, 41, 241
- Draine et al. (2014) Draine, B. T., Aniano, G., Krause, O., et al. 2014, ApJ, 780, 172
- Dudzevičiūtė et al. (2020) Dudzevičiūtė, U., Smail, I., Swinbank, A. M., et al. 2020, MNRAS, 494, 3828
- Dunne et al. (2022) Dunne, L., Maddox, S. J., Papadopoulos, P. P., Ivison, R. J., & Gomez, H. L. 2022, MNRAS, 517, 962
- Elbaz et al. (2018) Elbaz, D., Leiton, R., Nagar, N., et al. 2018, A&A, 616, A110
- Finkelstein et al. (2015) Finkelstein, S. L., Ryan, Russell E., J., Papovich, C., et al. 2015, ApJ, 810, 71
- Fudamoto et al. (2020) Fudamoto, Y., Oesch, P. A., Faisst, A., et al. 2020, A&A, 643, A4
- Fujimoto et al. (2022) Fujimoto, S., Brammer, G. B., Watson, D., et al. 2022, Nature, 604, 261
- Fujimoto et al. (2024) Fujimoto, S., Kohno, K., Ouchi, M., et al. 2024, ApJS, 275, 36
- Gentile et al. (2024) Gentile, F., Casey, C. M., Akins, H. B., et al. 2024, ApJ, 973, L2
- Gkogkou et al. (2023) Gkogkou, A., Béthermin, M., Lagache, G., et al. 2023, A&A, 670, A16
- Gottumukkala et al. (2024) Gottumukkala, R., Barrufet, L., Oesch, P. A., et al. 2024, MNRAS, 530, 966
- Greene et al. (2024) Greene, J. E., Labbe, I., Goulding, A. D., et al. 2024, ApJ, 964, 39
- Greve et al. (2008) Greve, T. R., Pope, A., Scott, D., et al. 2008, MNRAS, 389, 1489
- Gruppioni et al. (2020) Gruppioni, C., Béthermin, M., Loiacono, F., et al. 2020, A&A, 643, A8
- Guaita et al. (2022) Guaita, L., Aravena, M., Gurung-Lopez, S., et al. 2022, A&A, 660, A137
- Haider et al. (2016) Haider, M., Steinhauser, D., Vogelsberger, M., et al. 2016, MNRAS, 457, 3024
- Helton et al. (2024a) Helton, J. M., Sun, F., Woodrum, C., et al. 2024a, ApJ, 962, 124
- Helton et al. (2024b) Helton, J. M., Sun, F., Woodrum, C., et al. 2024b, ApJ, 974, 41
- Herard-Demanche et al. (2025) Herard-Demanche, T., Bouwens, R. J., Oesch, P. A., et al. 2025, MNRAS, 537, 788
- Hughes et al. (1998) Hughes, D. H., Serjeant, S., Dunlop, J., et al. 1998, Nature, 394, 241
- Jiang et al. (2018) Jiang, L., Wu, J., Bian, F., et al. 2018, Nature Astronomy, 2, 962
- Jiménez-Andrade et al. (2024) Jiménez-Andrade, E. F., Murphy, E. J., Momjian, E., et al. 2024, ApJ, 972, 89
- Kakimoto et al. (2024) Kakimoto, T., Tanaka, M., Onodera, M., et al. 2024, ApJ, 963, 49
- Kashino et al. (2023) Kashino, D., Lilly, S. J., Matthee, J., et al. 2023, ApJ, 950, 66
- Khusanova et al. (2021) Khusanova, Y., Bethermin, M., Le Fèvre, O., et al. 2021, A&A, 649, A152
- Kinney et al. (1996) Kinney, A. L., Calzetti, D., Bohlin, R. C., et al. 1996, ApJ, 467, 38
- Kodra et al. (2023) Kodra, D., Andrews, B. H., Newman, J. A., et al. 2023, ApJ, 942, 36
- Labbé et al. (2023) Labbé, I., van Dokkum, P., Nelson, E., et al. 2023, Nature, 616, 266
- Leitherer et al. (1999) Leitherer, C., Schaerer, D., Goldader, J. D., et al. 1999, ApJS, 123, 3
- Lemaux et al. (2022) Lemaux, B. C., Cucciati, O., Le Fèvre, O., et al. 2022, A&A, 662, A33
- Lestrade et al. (2022) Lestrade, J. F., Désert, F. X., Lagache, G., et al. 2022, A&A, 667, A23
- Li et al. (2024a) Li, Q., Conselice, C. J., Sarron, F., et al. 2024a, arXiv e-prints, arXiv:2405.17359
- Li et al. (2019) Li, Q., Narayanan, D., & Davé, R. 2019, MNRAS, 490, 1425
- Li et al. (2024b) Li, Z., Dekel, A., Sarkar, K. C., et al. 2024b, A&A, 690, A108
- Lim et al. (2024) Lim, S., Tacchella, S., Schaye, J., et al. 2024, MNRAS, 532, 4551
- Lin et al. (2025) Lin, X., Egami, E., Sun, F., et al. 2025, arXiv e-prints, arXiv:2504.08028
- Liu et al. (2018) Liu, D., Daddi, E., Dickinson, M., et al. 2018, ApJ, 853, 172
- Madau & Dickinson (2014) Madau, P. & Dickinson, M. 2014, ARA&A, 52, 415
- Matthee et al. (2024) Matthee, J., Naidu, R. P., Brammer, G., et al. 2024, ApJ, 963, 129
- McKinney et al. (2025) McKinney, J., Casey, C. M., Long, A. S., et al. 2025, ApJ, 979, 229
- McKinney et al. (2023) McKinney, J., Manning, S. M., Cooper, O. R., et al. 2023, ApJ, 956, 72
- Morishita et al. (2025) Morishita, T., Liu, Z., Stiavelli, M., et al. 2025, ApJ, 982, 153
- Morokuma-Matsui et al. (2025) Morokuma-Matsui, K., Yajima, H., & Abe, M. 2025, MNRAS, 539, 1834
- Narayanan et al. (2015) Narayanan, D., Turk, M., Feldmann, R., et al. 2015, Nature, 525, 496
- Neri et al. (2014) Neri, R., Downes, D., Cox, P., & Walter, F. 2014, A&A, 562, A35
- Noll et al. (2009) Noll, S., Burgarella, D., Giovannoli, E., et al. 2009, A&A, 507, 1793
- Oesch et al. (2023) Oesch, P. A., Brammer, G., Naidu, R. P., et al. 2023, MNRAS, 525, 2864
- Owen (2018) Owen, F. N. 2018, ApJS, 235, 34
- Pacifici et al. (2023) Pacifici, C., Iyer, K. G., Mobasher, B., et al. 2023, ApJ, 944, 141
- Perera et al. (2008) Perera, T. A., Chapin, E. L., Austermann, J. E., et al. 2008, MNRAS, 391, 1227
- Pérez-González et al. (2024) Pérez-González, P. G., Barro, G., Rieke, G. H., et al. 2024, ApJ, 968, 4
- Planck Collaboration et al. (2020) Planck Collaboration, Aghanim, N., Akrami, Y., et al. 2020, A&A, 641, A6
- Ponthieu et al. (2025) Ponthieu, N., Désert, F.-X., Beelen, A., et al. 2025, sub. to A&A
- Pope et al. (2005) Pope, A., Borys, C., Scott, D., et al. 2005, MNRAS, 358, 149
- Popesso et al. (2023) Popesso, P., Concas, A., Cresci, G., et al. 2023, MNRAS, 519, 1526
- Riechers et al. (2020) Riechers, D. A., Hodge, J. A., Pavesi, R., et al. 2020, ApJ, 895, 81
- Rodighiero et al. (2023) Rodighiero, G., Bisigello, L., Iani, E., et al. 2023, MNRAS, 518, L19
- Sommovigo et al. (2022) Sommovigo, L., Ferrara, A., Carniani, S., et al. 2022, MNRAS, 517, 5930
- Speagle et al. (2014) Speagle, J. S., Steinhardt, C. L., Capak, P. L., & Silverman, J. D. 2014, ApJS, 214, 15
- Sun et al. (2024) Sun, F., Helton, J. M., Egami, E., et al. 2024, ApJ, 961, 69
- Tacconi et al. (2020) Tacconi, L. J., Genzel, R., & Sternberg, A. 2020, ARA&A, 58, 157
- Tanaka et al. (2024) Tanaka, M., Onodera, M., Shimakawa, R., et al. 2024, ApJ, 970, 59
- Traina et al. (2024) Traina, A., Gruppioni, C., Delvecchio, I., et al. 2024, A&A, 681, A118
- Triani et al. (2020) Triani, D. P., Sinha, M., Croton, D. J., Pacifici, C., & Dwek, E. 2020, MNRAS, 493, 2490
- Walter et al. (2012) Walter, F., Decarli, R., Carilli, C., et al. 2012, Nature, 486, 233
- Wang et al. (2023) Wang, F., Yang, J., Hennawi, J. F., et al. 2023, ApJ, 951, L4
- Wang et al. (2019) Wang, T., Schreiber, C., Elbaz, D., et al. 2019, Nature, 572, 211
- Wang et al. (2024) Wang, T., Sun, H., Zhou, L., et al. 2024, arXiv e-prints, arXiv:2403.02399
- Weibel et al. (2024) Weibel, A., Oesch, P. A., Barrufet, L., et al. 2024, MNRAS, 533, 1808
- Witstok et al. (2023) Witstok, J., Jones, G. C., Maiolino, R., Smit, R., & Schneider, R. 2023, MNRAS, 523, 3119
- Xiao et al. (2024) Xiao, M., Oesch, P. A., Elbaz, D., et al. 2024, Nature, 635, 311
- Xiao et al. (2023) Xiao, M. Y., Elbaz, D., Gómez-Guijarro, C., et al. 2023, A&A, 672, A18
- Zavala et al. (2023) Zavala, J. A., Buat, V., Casey, C. M., et al. 2023, ApJ, 943, L9
- Zavala et al. (2021) Zavala, J. A., Casey, C. M., Manning, S. M., et al. 2021, ApJ, 909, 165
- Zavala et al. (2022) Zavala, J. A., Casey, C. M., Spilker, J., et al. 2022, ApJ, 933, 242
Appendix A Dust production
Observations have shown a relation between dust mass and stellar mass across an extensive redshift range. At all redshifts, the dust masses increase linearly with stellar mass. However, at fixed stellar mass, there is a large scatter of dust masses around the median, revealing the existence of a dust-poor galaxy population across all stellar masses (Triani et al. 2020). We compare on Fig. 5 our stellar-mass estimates with our inferred dust masses for the ten DSFG. We also show the stellar mass–dust mass relations derived by Witstok et al. (2023), based on dust production from stellar sources (including supernovae and AGB stars ; with and without reverse shock destruction which can lead to 95% dust destruction). Additionally, we include the relation corresponding to the highest possible metal yield (i.e. assuming all metals are locked up in dust grains), representing the theoretical maximum dust mass. Note that the exact normalization of these relations remains uncertain due to systematic effects, including the choice of IMF used to compute dust yields for a given stellar mass, and the range of plausible dust absorption cross-sections.
The dust masses observed in the majority of our galaxies (all but two) point to efficient dust formation. Indeed most of our galaxies are consistent with the dust masses expected from a simple, IMF-averaged yield of stellar sources alone (i.e., SN and AGB stars), assuming no dust destruction in the ISM or ejection from the galaxy. In general, it is challenging to populate our observed M⋆–Mdust relation using cosmological simulations, as our dusty galaxies occupy the upper envelope in both stellar and dust mass. They are, for instance, only sparsely represented in Di Cesare et al. (2023) and entirely absent in the sample of Triani et al. (2020).
The two most massive galaxies in our sample fall into regions where reverse-shock destruction of dust becomes significant, implying that additional dust growth mechanisms, such as grain growth in the interstellar medium, may be necessary to account for their observed dust masses. However, we caution that these two sources currently lack spectroscopic redshifts, and their stellar masses based on photometric estimates may be significantly overestimated. Robust spectroscopic confirmation is essential before drawing firm conclusions about their dust production.
For comparison, we also present the complete sample of N2CLS galaxies at from Berta et al. (2025), which occupy the same region i.e. near the SN+AGB dust production line. Notably, three galaxies lie above the maximum dust mass allowed: N2GN_1_17a (=3.19) and N2GN_1_32 (=3.65; detected in the Chandra 2 Ms map and requiring an AGN-torus component in the IRAC and MIPS bands) remain consistent within 1.3 (in Mdust), while N2GN_1_04 (=4.056, also known as GN20) is significantly offset.
The distribution of our measurements in the Mdust-M⋆ plane appears relatively flat. Sommovigo et al. (2022) showed that the Mdust–M⋆ relation differs between REBELS galaxies and ALPINE galaxies (even when using consistent [CII]-based methods to estimate Mdust), with REBELS galaxies exhibiting a flat trend, while ALPINE galaxies, as well as theoretical models (see Fig. 5 and their Fig. 3), suggest that more massive galaxies are dustier. Both samples probe lower dust and stellar masses than our NIKA2 sources. The observed flatness of the Mdust–M⋆ relation could partly arise from observational biases: the REBELS survey covers a limited stellar mass range, while for NIKA2, the 1.2 mm flux selection imposes a dust mass detection threshold.
@@ Appendix B Photometric redshifts with New-Hyperz
For the five sources with literature-based photometric redshifts at 5.2 (see B25 and references therein), we re-derived their redshifts independently using the updated photometry and the New-Hyperz SED-fitting code. New-Hyperz is a modified version of the public code Hyperz originally developped by Bolzonella et al. (2000). The library of templates used for this need includes: i) Eight evolutionary synthetic SEDs computed with the Bruzual & Charlot code (Bruzual & Charlot 2003), with Chabrier IMF (Chabrier 2003) and solar metallicity, matching the observed colors of local galaxies from E to Im types (namely a delta burst Single Stellar Population, a constant star-forming system, and six -models with exponentially decaying SFR); ii) Six empirical templates for galaxies in the local Universe: four SEDs compiled by Coleman et al. (1980), and two starburst galaxies from the Kinney et al. 1996 library; iii) Ten Starbursts99 evolutionary models (Leitherer et al. 1999), including e-lines, for single bursts and constant SFR models, each spanning five metallicities (Z 0.04, 0.02, 0.008, 0.004 and 0.001). The explored redshift domain goes from z0 to 16, without any prior in luminosity. The only restriction is that the age of the stellar population should be smaller than the age of the universe for a given redshift. Internal dust extinction is a free parameter that is allowed to take values between A0 and 6 magnitudes following the Calzetti et al. (2000) extinction law. During the SED-fitting process, to account in a realistic way for calibration uncertainties related to the combination of data coming from different facilities, across a wide range in wavelengths, we imposed a minimum photometric error of 0.2 magnitudes which is quadratically added to the original error bars. Indeed, in our experience this is needed to prevent artificially small error bars leading to low-quality fits, by correcting from effects such as biases between different instruments due to (absolute) zero-point uncertainties, photometry computed with an effectively different spatial sampling and spatial resolution, and contribution to the total flux coming from different spatial regions.
@@
The quality of the photometric redshifts with New-Hyperz and the settings described above for the SED-fitting has been checked using the 29 N2CLS galaxies with spectroscopic redshifts (Sect. 2), excluding N2GN_1_61 because it is only detected in one filter. Spectroscopic redshifts are in the range between z0.6 and 7.2. The results are as follows, with : = 0.058, the median of = -0.016, and the normalized median absolute deviation, which is less sensitive to outliers: = 1.48 median = 0.018. Regarding the fraction of outliers, defined in a conservative way as sources with , it is found to be 21% (6 objects/28), and only two galaxies (7%) exhibit . Among these six outliers, four exhibit a degenerate solution with a secondary solution consistent with . One of them is an X-ray AGN at zspec = 3.65 with a NIR-MIR SED dominated by a torus component, exhibiting a degenerate solution between z and z. Note that a poor zphot estimate is obtained for an obscured quasar at zspec=7.19 (Fujimoto et al. 2022).
Appendix C Contribution to the cosmic SFRD
Based on the five N2CLS galaxies with spectroscopic redshifts, we derive a cosmic star formation rate density (SFRD) of M⊙ yr-1 Mpc-3. Including the additional contribution from sources with photometric redshifts raises the total SFRD to M⊙ yr-1 Mpc-3. The quoted uncertainties combine Poisson errors and uncertainties on individual SFR estimates in quadrature. Figure 6 shows the contribution of the N2CLS galaxies within the overdensity at z to the cosmic SFRD. For comparison, we also display the total SFRD corrected for dust attenuation as compiled by Madau & Dickinson (2014). This comparison is discussed in the main text (Sect. 7.1). Additionally, we include the dust-obscured and unobscured components of the SFRD from Zavala et al. (2021). In their analysis, the contribution from dust-obscured galaxies peaks at –3 and declines at higher redshifts, reaching levels comparable to the unobscured star formation traced by rest-frame UV/optical surveys by . We find that the contribution of the dusty galaxies in the z overdensity is more than five times higher than the dust-obscured SFRD reported by Zavala et al. (2021) at similar redshifts.
Appendix D Notes about individual sources and JWST cutouts

### v2

@@ Acknowledgements.
We acknowledge financial support from the Programme National de Cosmologie and Galaxies (PNCG) funded by CNRS/INSU-IN2P3-INP, CEA and CNES, France, and from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (project CONCERTO, grant agreement No 788212). This work is based on observations carried out under project numbers 192-16 with the IRAM 30-m telescope, and projects W21CV, W23CX, and S24CF with NOEMA. IRAM is supported by INSU/CNRS (France), MPG (Germany) and IGN (Spain). This work is also based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. The observations are associated with programs #1181, 1895 and 3577. Some of the data products used here were retrieved from the Dawn JWST Archive (DJA). DJA is an initiative of the Cosmic Dawn Center (DAWN), which is funded by the Danish National Research Foundation under grant DNRF140. We thank Eric F. Jiménez-Andrade for sharing their data products from the 10 GHz Karl G. Jansky Very Large Array survey of GOODS-N. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. GL thanks Joris Witstok and Thomas Henning for enlightening discussions on dust production, and warmly thanks Fengwu Sun for sharing his Python code used to generate Figure 4.Appendix A Dust production
Observations have shown a relation between dust mass and stellar mass across an extensive redshift range. At all redshifts, the dust masses increase linearly with stellar masses. However, at fixed stellar mass, there is a large scatter of dust masses around the median, revealing the existence of a dust-poor galaxy population across all stellar masses (Triani2020). We compare in Fig. 6 our stellar-mass estimates with our inferred dust masses for the eight DSFG. We also show the stellar mass–dust mass relations derived by witstok2023 based on dust production from stellar sources (including supernovae and AGB stars; with and without reverse shock destruction, which can lead to 95% dust destruction). Additionally, we include the relation corresponding to the highest possible metal yield (i.e. assuming all metals are locked up in dust grains), representing the theoretical maximum dust mass. Note that the exact normalisation of these relations remains uncertain due to systematic effects, including the choice of IMF used to compute dust yields for a given stellar mass, and the range of plausible dust absorption cross-sections.
The dust masses observed in our galaxies point to efficient dust formation. Indeed our galaxies are consistent with the dust masses expected from a simple, IMF-averaged yield of stellar sources alone (i.e. supernovae and AGB stars), assuming no dust destruction in the ISM or ejection from the galaxy. In general, it is challenging to populate our observed M⋆–Mdust relation using cosmological simulations, as our dusty galaxies occupy the upper envelope in both stellar and dust mass. They are, for instance, only sparsely represented in 2023MNRAS.519.4632D and entirely absent in the sample of Triani2020.
We also show in Fig. 6 the location of N2GN_1_20, the most massive galaxy in our z5.2 sample, which is discussed in Appendix D. It falls into the region where reverse-shock destruction of dust becomes significant, implying that additional dust growth mechanisms, such as grain growth in the ISM, may be necessary to account for their observed dust masses. However, we caution that this source currently lack spectroscopic redshift, and robust spectroscopic confirmation is essential before drawing firm conclusions about its dust production.
For comparison, we also present the complete sample of N2CLS galaxies at from berta2025, which occupy the same region i.e. near the supernova+AGB dust production line. Notably, three galaxies lie above the maximum dust mass allowed: N2GN_1_17a (=3.19) and N2GN_1_32 (=3.65; detected in the Chandra 2 Ms map and requiring an AGN-torus component in the IRAC and MIPS bands) remain consistent within 1.3 (in Mdust), while N2GN_1_04 (=4.056, also known as GN20) is significantly offset.
The distribution of our measurements in the Mdust–M⋆ plane appears relatively flat. Sommovigo2022 show that the Mdust–M⋆ relation differs between REBELS galaxies and ALPINE galaxies (even when using consistent [CII]-based methods to estimate Mdust), with REBELS galaxies exhibiting a flat trend, while ALPINE galaxies, as well as theoretical models (see Fig. 6 and their Fig. 3), suggest that more massive galaxies are dustier. Both samples probe lower dust and stellar masses than our NIKA2 sources. The observed flatness of the Mdust–M⋆ relation could partly arise from observational biases: the REBELS survey covers a limited stellar mass range, while for NIKA2, the 1.2 mm flux selection imposes a dust mass detection threshold.
@@ Appendix B Photometric redshifts with New-Hyperz
For the four sources with literature-based photometric redshifts at (N2GN_1_20, N2GN_1_26, N2GN_1_43, N2GN_1_57), we re-derived their redshifts independently using the updated photometry and the New-Hyperz SED-fitting code. New-Hyperz is a modified version of the public code Hyperz originally developed by Bolzonella_2000. The library of templates used for this need includes: (i) Eight evolutionary synthetic SEDs computed with the Bruzual & Charlot code (bc03), with Chabrier IMF (chabrier2003) and solar metallicity, matching the observed colours of local galaxies from E to Im types (namely a delta burst single stellar population, a constant star-forming system, and six -models with exponentially decaying SFR); (ii) Six empirical templates for galaxies in the local Universe: four SEDs compiled by coleman1980, and two starburst galaxies from the Kinney1996 library; (iii) Ten Starbursts99 evolutionary models (STARBURST99), including e-lines, for single bursts and constant SFR models, each spanning five metallicities (Z 0.04, 0.02, 0.008, 0.004 and 0.001). The explored redshift domain goes from z0 to 16, without any prior in luminosity. The only restriction is that the age of the stellar population should be smaller than the age of the Universe for a given redshift. The internal dust extinction is a free parameter that is allowed to take values between A0 and 6 magnitudes following the Calzetti2000 extinction law. During the SED-fitting process, to account in a realistic way for calibration uncertainties related to the combination of data coming from different facilities, across a wide range in wavelengths, we imposed a minimum photometric error of 0.2 magnitudes, which is quadratically added to the original error bars. Indeed, in our experience this is needed to prevent artificially small error bars leading to low-quality fits, by correcting from effects such as biases between different instruments due to (absolute) zero-point uncertainties, photometry computed with an effectively different spatial sampling and spatial resolution, and contribution to the total flux coming from different spatial regions.
@@
The quality of the photometric redshifts with New-Hyperz and the settings described above for the SED-fitting has been checked using the 29 N2CLS galaxies with spectroscopic redshifts (Sect. 2), excluding N2GN_1_61 because it is only detected in one filter. Spectroscopic redshifts are in the range between z0.6 and 7.2. The results are as follows, with : = 0.058, the median of = -0.016, and the normalised median absolute deviation, which is less sensitive to outliers: = 1.48 median = 0.018. Regarding the fraction of outliers, defined in a conservative way as sources with , it is found to be 21% (6 of 28 objects), and only two galaxies (7%) exhibit . Among these six outliers, four exhibit a degenerate solution with a secondary solution consistent with . One of them is an X-ray AGN at zspec = 3.65 with a near-IR to mid-IR SED dominated by a torus component, exhibiting a degenerate solution between z and z. Also, a poor zphot estimate is obtained for an obscured quasar at zspec=7.19 (Fujimoto2022).
Appendix C Notes about individual sources and JWST cutouts